OpenStudy (anonymous):

LOG5 (-1/X) - LOG5 (-X) = -2 I'm working with log problems and have gotten down to the part where 1/5^2 = 1/x^2 but I either did something wrong or forgot a rule, possibly both :/ not very familiar with them

hartnn (hartnn):

why do you think thats wrong ?

hartnn (hartnn):

1/x^2 = 1/5^2 gives x^2 = 5^2 x =+5 or -5

hartnn (hartnn):

and x =+5 is not possible, so its x =-5

OpenStudy (anonymous):

Could you please show me the step where you are setting it equal to 0? and therefore get x= +or -5? I tend to forget the basics

hartnn (hartnn):

oh, this step : \(x^2 = 5^2 \\ x^2-5^2 = 0 \\ (x+5)(x-5) =0 \\ x+5 =0 ~ ~ OR ~~ x-5 =0 \\ x= -5, 5\)

OpenStudy (anonymous):

Yes that step thank you! :)

hartnn (hartnn):

welcome ^_^